OpenStudy (anonymous):

help please! show steps please idk how to do it: 4a^2+12ab+9b^2-25c^2

OpenStudy (anonymous):

a hint for u What is the numerator in the negative solution l2x+3l-6=14 Thank You!! Btw those are supposed to be absolute value signs The negative solution of l2x+3l-6=14 is: -(2x + 3) - 6 = 14 -2x - 3 - 6 = 14 -2x = 23 x = -23/2 The numerator of the negative solution is